Solve the following equation using the substitution method.
x^4 –5x^2 –14=0
Complete the following steps to solve the above equation:
Let u=x^2, substitute the variables, and write the new equation.
Now factor the new equation, apply the zero-product principle, and solve for u.
Now substitute u=x^2 to solve for x.

Answer by stanbon(75887)

Solve the following equation using the substitution method.
x^4 –5x^2 –14=0
----
Let u = x^2; then substitute:
----
u^2 -5u -14 = 0
------
Factor:
(u-7)(u+2) = 0
----
u = 7 or u = -2
------
Solve for "x":
x^2 = 7 or x^2 = -2
-----
x = +-sqrt(7) or x = +-i*sqrt(2)
=================================
Cheers,
Stan H.
=====================